OpenStudy (anonymous):

Ships A and B leave port together. For the next two hours, ship A travels at 10 mph in a direction 25° west of north while the ship B travels 25° east of north at 40 mph. (a) What is the distance between the two ships two hours after they depart? (answer in miles) (b) What is the speed of ship A as seen by ship B? (answer in mph)

OpenStudy (jamesj):

Right down an equation for each ship showing its position in Cartesian coordinates: r1(t) = (x1(t), y1(t)) r2(t) = (x2(t), y2(t)) Then the distance between them at t = 2 hours will be: | r1(2) - r2(2) | = sqrt( (x2(2) - x1(2))^2 + (y2(2) - y1(2))^2 ) and their relative velocity r2'(t) - r1'(t) or vice versa where ' = d/dt

OpenStudy (anonymous):